OpenStudy (anonymous):

Solve 4sin(2y - 0.3) + 5cos(27 - 0.3) = 0, for 0(lesser-than-or-equal-to) y (lesser-than-or-equal-to) "pi" radians.

OpenStudy (anonymous):

Tanvi is that 5cos(27-0.3)

OpenStudy (anonymous):

sorry, 2y - 0.3, not 27 -_-

OpenStudy (anonymous):

hhmm - i dont recall solving one like this before - i'll have to check my list of trig identities - wait a few minutes

OpenStudy (anonymous):

sure, suree, take your time. even if we replace (2y - 0.3) by a, then we would get 4sin(a) + 5cos(a) = 0. oh damn, i got it!!

OpenStudy (anonymous):

tats the way to go

OpenStudy (anonymous):

i don't know what that means, but it'll become tan (a) = -5/4, and then i'll solve it.

OpenStudy (anonymous):

yes - thats an easier way - do you really need anyone to help you tanvidals??!! lol well done

OpenStudy (anonymous):

yes, can you give me another value for a??? apart from -0.09 rad?? it needs to be positive.

OpenStudy (anonymous):

tan -5/4 - the possible values are between o and 2pi radians are in the second and 4th quadrant so we get pi - 0.9 and 2pi - 0.9

OpenStudy (anonymous):

i need between 0 and pi.

OpenStudy (anonymous):

yes so a = 2.24 , 5.38 so 2y - 0.3 = 2.24 y = 2.54 / 2 = 1.27 rad a=5.38 will be too big one value for y = 1.27 rad

OpenStudy (anonymous):

i needed to go higher than pi for value of a because y must have value 0 to pi

OpenStudy (anonymous):

thank youu :)

OpenStudy (anonymous):

yw - but i think ill have another coffee to stimulate my brain a bit!

OpenStudy (anonymous):

oh - plz click on good answer to close the question if you are happy with the answer

OpenStudy (anonymous):

i've just checked the answer on wolfram alpha - it gives it as 1.2727

OpenStudy (anonymous):

thats a good site for verification

OpenStudy (anonymous):

run your cursor over the root plot - the red spot on positive x-axis - i gives you this value